OpenStudy (loser66):

Find the Jacobian formula for f(x,y)= sin (x^2+y^3) Please, help. Formula only

OpenStudy (loser66):

@Empty

OpenStudy (loser66):

For some reason, I don't know how to write it down :( calculating is easy. :)

OpenStudy (loser66):

Can I do: let u = sin x, \[\dfrac{\partial f}{\partial x}=\dfrac{\partial }{\partial u}(sin (x^2+y^2))\dfrac{\partial (x^2+y^3)}{\partial x} \] \[\dfrac{\partial f}{\partial y}=\dfrac{\partial }{\partial u}(sin (x^2+y^2))\dfrac{\partial (x^2+y^3)}{\partial y} \]

OpenStudy (empty):

Isn't there another function, like g(x,y) that goes with this?

OpenStudy (loser66):

Nope, that is all I have. ha!! I have a bunch of them. I need know how to jot the formula down.

OpenStudy (irishboy123):

\[\frac{∂(x,y,z)}{∂(u,v,w)} =\det \left[\begin{matrix}\frac{∂x}{∂u} & \frac{∂y}{∂u} & \frac{∂z}{∂u} \\ \frac{∂x}{∂v} & \frac{∂y}{∂v} & \frac{∂z}{∂v}\\ \frac {∂x}{∂w} & \frac {∂z}{∂w} & \frac {∂z}{∂w}\end{matrix}\right]\] \[dxdydz = \left| \frac{∂(x,y,z)}{∂(u,v,w)} \right|dudvdw\]

OpenStudy (empty):

Jacobians are generally used to change coordinates, so this doesn't really make any sense, unless you're shifting from 2 dimensions onto 1 dimension which is kind of uncommon. Generally you have f(x,y) and g(x,y) which is the change from x and y to f and g while you can also look at x(f,g) and y(f,g) as two Jacobian matrices. The cool thing about these is that when you take all the partials and put them into their own matrices, they are matrix inverses of each other.

OpenStudy (empty):

So specifically: \[\left[\begin{matrix}f_x & f_y \\ g_x & g_y\end{matrix}\right]\left[\begin{matrix}x_f & x_g\\ y_f & y_g\end{matrix}\right] = \left[\begin{matrix}1 & 0\\ 0 & 1\end{matrix}\right]\]

OpenStudy (loser66):

This problem is \(f:\mathbb R^2\rightarrow \mathbb R\). It's qualify to find the Jacobian transformation.

OpenStudy (loser66):

\(f: (x,y)\mapsto sin (x^2+y^2)\)

OpenStudy (loser66):

*y^3

OpenStudy (empty):

I think you already have your answer, here, why isn't this enough? \[\dfrac{\partial f}{\partial x}=\dfrac{\partial }{\partial u}(sin (x^2+y^2))\dfrac{\partial (x^2+y^3)}{\partial x} \] \[\dfrac{\partial f}{\partial y}=\dfrac{\partial }{\partial u}(sin (x^2+y^2))\dfrac{\partial (x^2+y^3)}{\partial y} \]

OpenStudy (loser66):

One more question: what is difference between Derivative matrix and Jacobian ?

OpenStudy (empty):

Nothing, they are the same thing

OpenStudy (empty):

So for your case you have a 2x1 or 1x2 matrix of partial derivatives since you're going from 2 to 1 dimension.

OpenStudy (loser66):

Another one: :) what is difference between \(\dfrac{\partial }{\partial u}(sin (x^2+y^2))\)and \(\dfrac{\partial (sin (x^2+y^3))}{\partial x}\)

OpenStudy (empty):

Well partial u doesn't really make sense, it's just sort of like the chain rule. I see your mistake now I was tired and didn't pay attention but what you should have written to help you out with the chain rule is this: \[\large \text{let } u = x^2+y^2\]\[\large \frac{\partial}{\partial x} \sin(x^2+y^2)=\frac{\partial}{\partial x} \sin(u) \\ \large \cos(u)* \frac{\partial u}{\partial x}\] Now looking at what u is above we can evaluate \[\large \frac{\partial u}{\partial x} = \frac{\partial}{\partial x}(x^2+y^2)=2x\] So we plug it in and get \[\large \frac{\partial f}{\partial x} = 2x \cos(x^2+y^2)\]

OpenStudy (empty):

I guess I put the wrong power of y there, but it really doesn't matter much here haha.

OpenStudy (loser66):

Thanks for the help. I got it. :)

OpenStudy (empty):

Ok =)